# Reviewing Intelligent Cinematography: AI research for camera-based video production

### v1

@@ 3.1.3 Workflow Optimization and Automated Shot Composition
A semi-automated method introduced by Yu et al [285] was designed to handle object staging, automated camera positioning and shot composition, from an annotated script. The framework shown in Figure 8 optimizes camera parameters dependant on a generated sequence of actions for each present character. An action list is denoted by $\{a_{i}|i=1,2,...,N\}$, where $a_{i}$ is the $i$th action and $N$ is the total number of actions available. Following this, the action list is transformed to stage performance, where each action corresponds to a movement (time period) in a scene. This automates scene scheduling process. The consequent performance of an action within a scene is denoted by $\{p_{(t)_{a_{i}}},p_{(t+1)_{a_{i}}},...,p_{(t+l)_{a_{i}}}\}$, where $t$ denotes the moment the movement begins for $t\epsilon T$ and $(t+l)_{a_{i}}$, which denotes the sub-components of a movement for an action $a_{i}$ within a performance, $\{p_{t}|p_{1},p_{2},...,p_{T}\}$. During the camera optimization step, “aestethic” and “fidelity” models are jointly applied to a performance $p_{i}$, to determine the optimal camera $c_{t}$ to use at each time step $t$. The aesthetic model analyses six factors for camera planning: character visibility, character action, camera configuration, screen continuity (relative to character position), moving continuity (accounting for on-going changes between movements) and shot duration. The fidelity model first assumes that a mathematical model can approximate the relationship between a script and a generated video. To accomplish this, the model uses the global vector for word (GloVe) embedding model [186] to generate text from the generated video, then analyses the similarities between generated script and target script [^5] The final result is visualized here: https://www.youtube.com/watch?v=0PUdV6OeMac.. A GloVe embedding is vector representation for words that are trained using the global co-occurrence word-word of word pairs/groups. Thus, sub-structures of the vector space can define synonyms (proximal parallel vectors) and canonical structures (vector paths). This is similar to the latent variables for text-based generative AI, discussed in Section 2.2.
Figure 8: Illustration of the pipeline for text2animation in Yu et al [285], where $c_{j}$ is the camera selected for capturing action $a_{i}$.
@@
Overall, both of these methods produce naive approaches to workflow optimization which is constrained by cinematographic understanding and current technological capability. For example, the crux of the evaluation mechanism in Yu et al [285] depends not only on the assumption that a fidelity model can be constructed, but also the ability to re-translate a video into text in a manner appropriate for evaluation. The reliance on this chained processes reduces the effectiveness of this model. Though as Yu et al [285] concludes, there is still much to accomplish in optimising production workflows. For example, a further extension of method like GLAMOUR could involve automating the generation of annotated shots or for extending story-board process. This could prove beneficial for productions with tighter budgets and/or deadlines. Additionally, an improved method of evaluating such models is necessary, as we have doubts as to the validity of text-based testing for cinematographic models.
@@
Considering that workflows are often fragmented due to reliance on external collaborators, research focused on supporting collaboration seems more plausible. For example, a model could be tasked with acting as an administrator for a shared resource - deciding if committed changes should be accepted given some camera-based constrains. Leaning into ICVFX, cameras such as the Axibo PT4666Product information: https://www.axibo.com/product/pt4. motion controlled slider host a cloud-based Unreal Engine workflow and directly attend to the shared resource paradigm. Use of such tools could be extended to investigations projecting dynamic assets [65, 134, 116] and validating the in-camera (artistic) composition given a feature-based target [138, 265].
@@ 3.1.4 Challenges and Future Work
One of the challenges with introducing novel AI to the production process is testing and validation. Unfortunately, real production environments are not only challenging to access but difficult to control. Therefore most proposed research favours testing in simulation through virtual environments. Clearly, there is some disparity between a semi-informed simulation and a real use case. Going forward, we need to acknowledge the copious benefit of experimental productions [91]. Experimental productions could become a platform for testing IC tools, whether this is to accomplish menial tasks so artists can focus on creative experimentation or to explicitly support one’s creative expression. Alternatively, we could envision using 3-D reconstructions of real scenes, discussed in Section 3.2, to achieve more reliable tests on real scenarios. This would provide more relatable results for cinematographers that intend using IC to direct cameras in real environments, as well as those that wish capture cinematic footage using 3-D reconstructions.
@@
With respect to supporting the cinematographic language, there are many considerations to make. For example, the application of language can be beneficial for semantic script analysis/generation [101, 71, 161] but it could may also act as a form of communication between production staff and AI interfaces [55, 285]. We found PSL to be particularly interesting as it is represented as mid-level language. This is beneficial as it reduces the reliance of solving natural language processing (NLP) problems (e.g. a sequence2sequence encoding/decoding problem), which can be troublesome when confronted with complex shot descriptions or tasked with describing subjective cinematographic observations. Though, we should acknowledge the progress made in NLP and semantic analysis research over the last decade [34, 44]. Consequently, we are interested to see more proposals for learning linguistic structures for cinematographic production. For example, riding off recent breakthroughs in NLP [200, 165] researchers could explore LLMs for camera control and workflow optimization tasks. Consequently, we believe the progression of this will rely on establishing reliable sources of data. For example, for semantic shot analysis one could propose a comprehensive data set and method of labelling shots using PSL to reduce the complexity of natural language structures for word-embeddings.
@@
Otherwise, DA research provides a platform to host a range of AI controllers. They can be adapted for specific purpose and architectures are often implemented as modular pipelines. However, evaluating performance in-the-wild is limited by access to real sets, production staff and the limited number of benchmark datasets and models. This indicates that the state of DA research still has a considerable journey before solutions can be applied on a commercial scale. We hypothesize that as production is further facilitated in other areas, such as improved camera control [5, 57, 114], DA could be facilitated by controlling a set of automated tasks on a higher-plane of abstraction, perhaps through the use of semantic query languages for resource descriptive frameworks (RDF) [62, 89] or DCCL and PSL.
@@
With respect to workflow optimization strategies, we primarily question the social benefit of MP research. As discussed in Section 1, IC relies on easing the production of artistic film-making, so automating the entire process (‘text-to animation’) leaves little room for creative input [269, 192, 8]. However, this could be used to support the story-boarding process and other pre-production tasks. Taking this further, one could formulate a storyboard as a set of animated clips to make decisions on shot composition. Thus future work could look at optimising shot composition, such as for style transfer [265] or attention optimization [252, 188].
@@ 3.2 Virtual Production
In this section we focus on virtual production using real cameras. This is important as it delineates from classical virtual production used in animation or gaming, whereby two notable differences exist. Firstly, compositing virtual scenes for animation is less challenging than compositing virtual assets in a real scene for an IC production as nuances in lighting, colouring and perspective between real and virtual assets require attention. Secondly, real cameras are limited by physical and fiscal constraints, like set topology, additional hardware and technical skill for achieving specific camera motions. These matters are trivial to accomplish with a virtual camera that has 6DoF.
These differences underpin the general concerns of research tied to virtual production for IC. Hence, in this section we look at research involved with ICVFX and LED Volumes - touching upon works that investigate re-colourization and image-based lighting (IBL). We also discuss research focused on synthesizing virtual replicas of real actors and scenes through NeRFs, prefaced in Section 2.5. This looks at removing the physical and fiscal limitations of using real cameras for content acquisition, allowing users to re-capture real scenes in the context of a virtual environment/engine, as is achieved with classical virtual production.
@@ 3.2.1 ICVFX and LED Volumes
Numerous scholarly works have undertaken the task of dissecting visual effects (VFX) and in-camera applications within the realm of cinematography. However, the relevant use case(s) for IC is vaguely defined. Considering a purely cinematographic perspective, ICVFX usually involves CGI and/or compositing techniques that are executed in real time, providing cinematographers with a live feed of how virtual effects will appear relative to the real scene set-up. This can support production in numerous ways, such as an indicator for poor lighting or as a way to pre-visualize compositing and VFX to ease post production challenges. Here, the technical practice mainly concerns the ability to distinguish foreground and background elements [228] as well as modify lighting and colour [18]. Sharma [228] briefly presents chroma key and roto scoping as paradigms, with a focus on the former; in Figure 10(a). While, Bengtsson and Kang [18] presents how IBL, a well understood method in VFX in Figure 9, has evolved into a lighting workflow surrounding LED video screens for driving in-camera relighting and colourization, Figure 10(b). This hints at two distinct use cases for IC research:
@@
    Automatic IBL and re-colourization
Figure 9: IBL maps an environment image onto a 3-D primitive (e.g. a sphere). The IBL environment is treated as a light source and using ray tracing a model is composited.
@@
Figure 10: Illustrating ICVFX compositing approaches
@@
It is generally known that use case (1) is readily accomplished by using chroma screens (1.i) and keying out the colour corresponding to the screen. However this introduces issues for colour-based image segmentation. Cheng et al [50] touches upon relevant colour-based image segmentation techniques angled towards the wider AI audience. The authors discuss solutions such as histogram thresholding [149], using binary trees to store data intensive 3-D colour spaces [225, 222], region-based methods [249, 49], fuzzy techniques [110, 247] and neural networks [106, 35]. Consequently the authors highlight problems with shading (e.g. shadows and highlights) and texturing. The most apparent case of this is interfering light-bounce from the chroma screen. This leads to subject specific approaches, for example Sigal et al [232] approaches real-time skin segmentation for video opposed by time-varying illumination. Sigal et al [232] optimizes a second order Markov model to predict a skin-colour histogram over time. Additionally, chroma keying restricts the ability to present colours similar to that of the chroma screen.
@@
More recently, we see approaches attending to case (1.ii). This is a popular paradigm outside of cinematography whereby we see emerging research such as the Segment Anything Model (SAM) [129] [^8] Still in pre-print. - a one-click solution to general image segmentation. Angled towards cinematography we find work such as Roto++ [142] - a rotoscoping tool with the ambition of respecting the artists’ requirements. Roto++ improves upon traditional interpolation techniques by combining a real-time appearance tracking and a novel shape manifold regularization process (built on the Gaussian process latent variable model (GP-LVM) [135]). Subsequently, within a sequence of frames the method (1) predicts the change of a shape manifold and (2) identifies which next keyframe needs to be manually roto scoped.
@@
Concerning case (2); while IBL is classically a 3-D rendering technique, it can be applied to real scenes where the lighting set-up can be readily changed [67]. Ren et al [209] and Wang et al [259] discuss approaches that reconstruct the light transport matrix (LTM) [^9] This defines light interactions on object surfaces.. Wang et al [259] classifies the approaches into three categories: (i) brute force [179, 92], directly modelling the LTM, (ii) sparsity based [84, 162, 205], modelling a set of basis functions under the assumption that each row of the LTM can be linearly approximated, and (iii) coherence based [181, 259, 79], analysing the coherence reflectance field to acquire the LTM. The limitation with these methods is that they require multiple images under varying lighting condition, meaning for video this problem is more challenging to address.
@@
Interestingly focus on cases (1) and (2) has shifted towards using LED panels to project virtual backgrounds as a practical solution. For case (1) foreground-background separation is made easy by replacing chroma screens with virtual backgrounds displayed live on interconnected LED panels. While more costly and energy intensive, it reduces the workload for post-production and avoids the need to roto scope. Unfortunately this means the need for research in this area is minimal and shifts toward supporting computation (e.g. reducing energy consumption). For case (2) the outcome is not so severe. Instead, the introduction of LED panels offers new possibilities for automated lighting calibration, now including the LED panels as a light source [174, 185, 112, 96]. For example, LeGendre et al [139] treats the panel lighting as ambient light, producing an example result in Figure 11
@@
Figure 11: Results from images taken in natural lighting (left) and configuring the LED panel for lighting (right) provided by LeGendre et al [139]
@@
[139] accomplishes this by first applying two matrices, $M$ and $N=MQ^{-1}$ to the out-of-camera-frustum and in-camera-frustum, respectively, where $M$ and $N$ represents a 3$\times$3 pre-correction matrix. Then applying the post-correction matrix, $Q$, to the final image, where $Q$ represents a 3$\times$3 post-correction matrix that re-maps viewed pixels to the desired/expected colour schema. $M$ is solved through matrix calculation, from known LED emission spectral sensitivity functions, i.e. $M=[SL]^{-1}$, where $[SL]$ represents the observed average pixel values from capturing light emitted by the LED panels. $Q$ is found by minimising the squared error between predicted pixel values and target pixel values, using the 3x3 matrix $[SRL]_{j}$. This encodes the spectral modulation and integrates the camera spectral sensitivity functions and LED emission and material reflectance spectra [139] for a given colour chart [^10] A colour chart is used to map shades and tints of red, green and blue, where each square in a chart is a different shade. square, $j$. Testing on colour charts showed near-optimal results, though resulting errors are limited by using a 3$\times$3 linear transformation kernel. LeGendre et al [139] highlights issues resulting in de-saturation of skin colour [^11] The authors propose further testing on a larger spectrum of skin colours. and fabrics; which is reasoned by the restrictive colour schema from LED panel lighting.
@@
Similarly, Smith and Zink [234] looks at correcting dynamic in-camera-frustum hue changes, though using simpler colour transforms. Additionally, Debevec and LeGendre [68] investigates a method for HDR-Image (HDRI) lighting reproduction, going from virtual HDRI setting to a LED volume setting. This essentially inverts the classical IBL problem and is approached by dilating pixels above a given threshold to meet constrains on local average pixel values displayed displayed on a virtual LED wall.
@@
Overall, the use of LED panelling in production is still new. Aside from discussing lighting and colourization, the IC landscape outside of this paradigm is undefined [125]. Subsequently, we have aggregated a set of non-scientific resources, in Table 1, which contribute to cinematographic discussions concerning the use of LED panels. We additionally categorized the list to provide further clarity on subject materials.
@@
Table 1: Supplementary Online Resources containing articles, documentation and video-based discussions.

### v3

@@ 3.1.3 Workflow Optimization and Automated Shot Composition
A semi-automated method introduced by Yu et al [298] was designed to handle object staging, automated camera positioning and shot composition, from an annotated script. The framework shown in Figure 8 optimizes camera parameters dependent on a generated sequence of actions for each present character. An action list is transformed to a stage performance, where each action corresponds to a movement (time period) in a scene. This automates the scene scheduling process. During the camera optimization step, “aesthetic” and “fidelity” models are jointly applied to a performance, to determine the optimal camera to use at each time step. The aesthetic model analyses six factors for camera planning: character visibility, character action, camera configuration, screen continuity (relative to character position), moving continuity (accounting for on-going changes between movements) and shot duration. The fidelity model first assumes that a mathematical model can approximate the relationship between a script and a generated video. To accomplish this, the model uses the global vector for word (GloVe) embedding model [196] to generate text from the generated video, then analyses the similarities between generated script and target script [^3] The final result is visualized here: https://www.youtube.com/watch?v=0PUdV6OeMac.. A GloVe embedding is vector representation for words that are trained using the global co-occurrence word-word of word pairs/groups. Thus, sub-structures of the vector space can define synonyms (proximal parallel vectors) and canonical structures (vector paths). This is similar to the latent variables for text-based generative AI, discussed in Appendix A.4.
Figure 8: Text2animation for shot selection discussed in Yu et al [298], where $c_{j}$ is the camera selected for capturing action $a_{i}$ during the performance $p_{k}$ and relies on jointly learning an aesthetic and fidelity model
@@
Overall, both of these methods produce naive approaches to workflow optimization - constrained both by cinematographic understanding and current technological capability. For example, the crux of the evaluation mechanism in Yu et al [298] depends not only on the assumption that a fidelity model can be constructed, but also the ability to re-translate a video into text in a manner appropriate for evaluation. The reliance on this chained processes reduces the effectiveness of this model. Though as Yu et al [298] concludes, there is still much to accomplish in optimizing production workflows. For example, a further extension of method like GLAMOUR could involve automating the generation of annotated shots or for extending story-board process. This could prove beneficial for productions with tighter budgets and/or deadlines. Additionally, an improved method of evaluating such models is needed, as we have doubts as to the validity of text-based testing for cinematographic models.
@@
Considering that workflows are often fragmented due to reliance on external collaborators, research focused on supporting collaboration seems appropriate. For example, a model could be tasked with acting as an administrator for a shared resource - deciding if committed changes should be accepted given some camera-based constrains. Leaning into ICVFX, cameras such as the Axibo PT4444Product information: https://www.axibo.com/product/pt4. motion controlled slider host a cloud-based Unreal Engine workflow and directly attend to the shared resource paradigm. Use of such tools could be extended to investigations projecting dynamic assets [67, 144, 123] and validating the in-camera (artistic) composition given a feature-based target [148, 277].
@@ 3.1.4 Challenges and Future Work
One of the challenges with introducing novel AI into the production process is how to test and validate it. Unfortunately, real production environments are not only challenging to access but also difficult to control. Therefore most proposals favor testing using simulation via virtual environments. Clearly, there is disparity between a semi-informed simulation and a real use case. Going forward, we need to acknowledge the copious benefit of experimental productions [97]. Experimental productions could become a platform for testing IC tools, whether this is to accomplish menial tasks so artists can focus on creative experimentation or to explicitly support one’s creative expression. Examples of this already exist, such as Lux Aeterna who produced a sci-fi movie, RENO, as part of the MyWorld Strength in Places Programme, as a means of exploring AI tool-sets in visual effects, Hifle [106]. This RENO production was used to explore the use of AI in post-production, however such work could be elevated to production-oriented tasks. Alternatively, we can envision using 3-D reconstructions of real scenes, discussed in Section 3.2, to achieve more reliable tests on real scenarios. This would provide more relatable results for cinematographers that intend using IC to direct cameras in real environments, as well as those that wish capture cinematic footage using 3-D reconstructions.
@@
With respect to supporting the cinematographic language, there are many considerations. For example, the application of language can be beneficial for semantic script analysis/generation [108, 73, 171] but it may also act as a form of communication between production staff and AI interfaces [57, 298]. We found PSL to be particularly interesting as it is represented as mid-level language. This is beneficial as it reduces the reliance of solving natural language processing (NLP) problems (e.g., a sequence2sequence encoding/decoding problem), which can be troublesome when confronted with complex shot descriptions or tasked with describing subjective cinematographic observations. However, we must acknowledge the progress made in NLP and semantic analysis research over the last decade [35, 46], which could lead to better proposals for learning linguistic structures for cinematographic production. For example, riding off recent breakthroughs in NLP [211, 175] researchers could explore LLMs for camera control and workflow optimization tasks, reliant of course on establishing reliable sources of data. For example, for semantic shot analysis one could propose a comprehensive dataset and method of labeling shots using PSL to reduce the complexity of natural language structures for word-embeddings.
@@
Otherwise, DA research provides a platform to exploit a range of AI controllers. These can be adapted for specific purposes and architectures are often implemented as modular pipelines. However, evaluating performance in-the-wild is limited by access to real sets, production staff and the limited number of benchmark datasets and models. This indicates that the state of DA research still has a considerable journey before solutions can be applied on a commercial scale. We hypothesize that as production is further facilitated in other areas, such as improved camera control [5, 59, 121], DA could be facilitated by controlling a set of automated tasks on a higher-plane of abstraction, perhaps through the use of semantic query languages for resource descriptive frameworks (RDF) [64, 95] or DCCL and PSL.
@@
With respect to workflow optimization strategies, we must examine the social benefit of MP research. As discussed in Section 1, the aim of IC is to better facilitate artistic film-making. However, automating the entire process (from ‘text-to-animation’) may leave little room for creative input [282, 202, 190]. MP could be used to support story-boarding and other pre-production tasks. Taking this further, one could formulate a storyboard as a set of animated clips to make decisions on shot composition. Thus future work could look at optimizing shot composition, such as for style transfer [277] or attention optimization [264, 198].
@@ 3.2 Virtual Production
In this section we focus on virtual production using real cameras. This is important as it delineates from classical virtual production, used in animation or gaming. Two notable differences exist: firstly, compositing virtual scenes for animation is less challenging than compositing virtual assets in a real scene for an IC production as nuances in lighting, coloring and perspective between real and virtual assets require attention. Secondly, real cameras are limited by physical and fiscal constraints, like set topology, additional hardware and technical skill for achieving specific camera motions. These matters are trivial to accomplish with a virtual camera that has 6 Degrees of Freedom (6DoF).
These differences underpin the challenges of research tied to virtual production for IC. Hence, in this section we review research relating to ICVFX and LED Volumes - touching upon works that investigate re-colorization and image-based lighting (IBL). We also discuss research focused on synthesizing virtual replicas of real actors and scenes through NeRFs, prefaced in Section 2.4. This looks at removing the physical and fiscal limitations of using real cameras for content acquisition, allowing users to re-capture real scenes in the context of a virtual environment/engine, as is achieved with classical virtual production.
@@ 3.2.1 ICVFX and LED Volumes
Numerous scholarly works have undertaken the task of dissecting visual effects (VFX) and in-camera applications within the realm of cinematography. However, the relevant use case(s) for IC is vaguely defined. Considering a purely cinematographic perspective, ICVFX usually involves CGI and/or compositing techniques that are executed in real time, providing cinematographers with a live feed of how virtual effects will appear relative to the real scene set-up. This can support production in numerous ways, such as providing an indicator for poor lighting or as a way to pre-visualize compositing and VFX to ease post-production challenges. Here, technical practice mainly involves distinguishing foreground and background elements [240] as well as modifying lighting and color [18]. Sharma [240] briefly presents chroma key and roto scoping as paradigms, with a focus on the former; as illustrated in Figure 10(a). While, Bengtsson and Kang [18] presents how IBL, a well understood method in VFX in Figure 9, has evolved into a lighting workflow surrounding LED video screens for driving in-camera relighting and colorization, Figure 10(b). This hints at two distinct use cases for IC research:
@@
    Automatic IBL and re-colorization
Figure 9: Image based lighting process for compositing: IBL maps an environment image onto a 3-D primitive (e.g., a sphere). The IBL environment is treated as a light source and using ray tracing a model is composited.
@@
Figure 10: Approaches to in-camera compositing: using automatic segmentation, chroma keying and LED stage/volume
@@
It is generally known that use case (1) is readily accomplished by using chroma screens (1.i) and keying out the color corresponding to the screen. However this introduces issues for color-based image segmentation. Cheng et al [52] touches upon relevant color-based image segmentation techniques angled towards the wider AI audience. The authors discuss solutions such as histogram thresholding [159], using binary trees to store data intensive 3-D color spaces [237, 234], region-based methods [261, 51], fuzzy techniques [118, 259] and neural networks [114, 36]. Consequently Cheng et al [52] highlight problems with shading (e.g., shadows and highlights) and texturing. The most apparent case of this is interfering light-bounce from the chroma screen. This leads to subject specific approaches, for example Sigal et al [244] approaches real-time skin segmentation for video opposed by time-varying illumination. Sigal et al [244] optimizes a second order Markov model to predict a skin-color histogram over time. Additionally, chroma keying restricts the ability to present colors similar to that of the chroma screen.
@@
More recently, we have observed approaches attending to case (1.ii). This is a popular paradigm outside of cinematography whereby we see emerging research such as the Segment Anything Model (SAM) [138] - a one-click solution to general image segmentation. Angled towards cinematography we find work such as Roto++ [152] - a rotoscoping tool with the ambition of respecting the artists’ requirements. Roto++ improves upon traditional interpolation techniques by combining a real-time appearance tracking and a novel shape manifold regularization process (built on the Gaussian process latent variable model (GP-LVM) [145]). Subsequently, within a sequence of frames the method (1) predicts the change of a shape manifold and (2) identifies which next keyframe needs to be manually roto scoped.
@@
Concerning case (2); while IBL is classically a 3-D rendering technique, it can be applied to real scenes where the lighting set-up can be readily changed [69]. Ren et al [220] and Wang et al [271] discuss approaches that reconstruct the light transport matrix (LTM) [^6] This defines light interactions on object surfaces.. Wang et al [271] classifies the approaches into three categories: (i) brute force [189, 98], directly modeling the LTM, (ii) sparsity based [89, 172, 216], modeling a set of basis functions under the assumption that each row of the LTM can be linearly approximated, and (iii) coherence based [191, 271, 84], analysing the coherence reflectance field to acquire the LTM. The limitation with these methods is that they require multiple images under varying lighting condition, meaning for video this problem is more challenging to address.
@@
Interestingly, focus on cases (1) and (2) has shifted towards using LED panels to project virtual backgrounds as a practical solution. For case (1) foreground-background separation is made easy by replacing chroma screens with virtual backgrounds displayed live on interconnected LED panels. While more costly and energy intensive, it reduces the workload for post-production and avoids the need to roto scope. Unfortunately this means the need for research in this area is minimal and shifts toward supporting computation (e.g., reducing energy consumption). For case (2) the outcome is not so severe. Instead, the introduction of LED panels offers new possibilities for automated lighting calibration, now including the LED panels as a light source [184, 195, 102]. For example, LeGendre et al [149] treats the panel lighting as ambient light, producing an example result in Figure 11
@@
Figure 11: Results from images taken in natural lighting (left) and configuring the surrounding LED panel stage/volume for in-camera lighting (right) provided by LeGendre et al [149]
@@
[149] accomplishes this by investigating a set of pre-correction (kernel) matrices applied to inner and outer LED wall frustums. Then with a separate post-correction matrix applied to the in-camera image, pixels are re-maped to the desired/expected color schema. The pre-correction matrices are solved by comparing and minimizing the difference between the expected LED emission spectral sensitivity functions and the resulting sensitivity function captured by the camera. This optimization is accomplished for a given color chart [^7] A color chart is used to map shades and tints of red, green and blue, where each square in a chart is a different shade.. Testing on a variety of color charts showed near-optimal results, though resulting errors are limited by the size of the pre- and post-correction kernels. LeGendre et al [149] highlights issues resulting in de-saturation of skin color [^8] The authors propose further testing on a larger spectrum of skin colors. and fabrics; which is reasoned by the restrictive color schema from LED panel lighting. We provide a deeper technical explanation of this method in Appendix A.3.
@@
Similarly, Smith and Zink [246] looks at correcting dynamic in-camera-frustum hue changes, though using simpler color transforms. Additionally, Debevec and LeGendre [70] investigates a method for HDR-Image (HDRI) lighting reproduction, going from virtual HDRI setting to a LED volume setting. This essentially inverts the classical IBL problem and is approached by dilating pixels above a given threshold to meet constrains on local average pixel values displayed displayed on a virtual LED wall.
@@
Overall, the use of LED paneling in production is still relatively new. Aside from discussing lighting and colorization, the IC landscape outside of this paradigm is undefined [134]. Subsequently, we have aggregated a set of non-scientific resources, in Table 1, which contribute to cinematographic discussions concerning the use of LED panels. We additionally categorized the list to provide further clarity on subject materials.
@@
Table 1: Online Resources containing articles, documentation and video-based discussions on virtual stage/volume film production.